### v1

@@ 4.3.1 Classification Metrics
ROC Curve and AUC (Area Under the Curve): The ROC curve is a graphical representation of a classification model’s performance across all classification thresholds. It plots the True Positive Rate (TPR) against the False Positive Rate (FPR). The AUC provides a single scalar value to summarize the model’s performance; a higher AUC indicates better model performance.
@@ 5.1 The Performance Evaluation of TabulaTime in ACS Prediction
The evaluation of the TabulaTime model was conducted through a series of experiments to assess its predictive performance for description of Acute Coronary Syndrome (ACS) risk. The experiments were designed to evaluate the influence of time-series air pollution data, compare TabulaTime with existing machine learning models, and analyse its generalization capabilities under varying conditions.
@@ 5.1.1 Impact of Time-Series Length for ACS Prediction
The model was tested with time-series air pollution data of varying lengths (3, 7, 10, and 14 days) to identify the optimal time window for ACS prediction. As shown in Table 2, the 10-day time series achieved the best results, with an accuracy of 82.3% and an AUC of 0.91, outperforming other time lengths across all metrics. This suggests that a 10-day window strikes a balance between capturing essential trends and avoiding noise introduced by excessive data. Shorter windows (3 and 7 days) lacked critical information, while longer windows (14 days) diluted key patterns.
@@ 5.1.2 Comparison With Existing Machine Learning Models for ACS Risk Prediction
Table 3: Model performance between traditional machine learning method and proposed method.
@@
Both models showed an improvement when air pollution is included as a factor. The performance of RF improves by 4.5% (from 0.6 to 0.627), LightGBM improved by 1.6% (from 0.645 to 0.655) and Catboost improves by 3.5% % (from 0.665 to 0.688). TabulaTime's performance improved by 10.1% (from 0.753 to 0.829). This indicated that incorporating air pollution data enhanced the predictive capabilities of both models, but the improvement was more pronounced for TabulaTime. Figure 5 and Figure 6 reported the Area Under the Receiver Operating Characteristic Curve (AUC-ROC) for the best-performing machine learning model (Catboost) and the proposed TabulaTime. The results confirm that the model performance improved with the inclusion of air pollution indicators. The ROC of the Catboost increased from 0.67 to 0.72, while the proposed TabulaTime improved the prediction accuracy from 0.83 to 0.91.
Figure 5: ROC curve of Catboost with and without air pollution.
Figure 6: ROC curve of TabulaTime with and without air pollution.
@@ 5.2.1 Classification
Table 4: Model performance on classification tasks.
@@ 5.2.2 Forecasting
Table 5: Model performance on forecasting tasks.
Figure 8: The forecasting result of PatchRWKV on Weather (a and b) and Salford Air Pollution dataset (c,d,e and f)
Table 5 presents the performance of various time series processing model on two forecasting tasks. The models are categorized into four types: MLP-based, CNN-based, Transformer-based, and RNN-based. Performance was evaluated using two metrics: Mean Squared Error (MSE) and Mean Absolute Error (MAE). The PatchRWKV approach demonstrates superior performance on both datasets, achieving the lowest MSE and highly competitive MAE values. Figure 8 showed the forecasting results of PatchRWKV on the two real datasets forecasting tasks. The model results were successful in predicting the trends of the pollution indicators. This indicates that PatchRWKV is highly effective in terms of both accuracy and robustness compared to other methods evaluated. Meanwhile, the RNN-based models generally outperformed other types in weather forecasting, with PatchRWKV and LSTM showing particularly strong results.
@@ 5.3 Feature Importance and Model Interpretability
Figure 9: Feature importance on ACS STEMI VS NSTEMI prediction.
In this work, we reported feature importance to explain how models make decisions to predict STEMI VS. NSTEMI.The analysis highlights the contributions of 25 risk factors (Figure 9), encompassing patient demographics, clinical history, vital signs, medications, diagnostic indicators, and air pollution metrics. The analysis revealed that Systolic Blood Pressure (SBP) was the most influential predictor of STEMI and NSTEMI, with the highest importance score (0.084), emphasizing its critical role in cardiovascular function. Body Mass Index (BMI) followed as the second most important feature (0.073), highlighting the metabolic influence on cardiovascular health, particularly in chronic conditions. Air quality metrics, such as PM10sum, PM10max, and NO2sum, were also significant contributors, underscoring the acute and cumulative impact of air pollution on cardiovascular risk. Moderately important features included creatinine levels (0.055), indicative of renal function, and age at admission (0.064), a well-established cardiovascular risk factor. Additionally, historical factors like angina and smoking status provided valuable insights into patient-specific risks. While metabolic factors like diabetes and hypertension were relevant, their lower importance scores suggests a more indirect role in the model. Lesser but notable contributors included peripheral vascular disease and prior CABG/PCI, reflecting the background conditions associated with ACS risk.
Figure 10: Day-wise Importance for ACS Prediction.
@@
Figure 10 showed the average importance of air qiality each day for predicting ACS risk. The daily step importance analysis revealed that pollution levels on the day prior to a clinical presentation of ACS (Day 1) held the highest predictive significance, followed by a gradual decline in relevance for preceding days. This finding emphasizes the acute impact of short-term pollution exposure on presentation of ACS.
@@
STEMI NSTEMI T-test p-value 25% 50% 75% 25% 50% 75% Systolic BP 114 128 146 121 138 156 0.0019 (<0.05) Creatinine 70.75 86 99 72 88 108 0.0049 (<0.05) Age At Admission 59 71 80.25 60 72 81 0.4911 (>0.05) BMI 23.36 26.02 29.5 24.53 28.03 32.23 0.0021 (<0.05) PM10sum 748.94 910.2 1226.46 769.85 978.45 1379.95 0.0487 (<0.05) NOsum 160.45 345.22 751.63 157.07 296.12 681.47 0.4883 (>0.05) NO2sum 1092.68 1502.32 2045.5 1112.78 1462.57 2049.09 0.628 (>0.05) NOXsum 1318.5 2060.17 3244.01 1388.16 1916.61 3054.62 0.5147 (>0.05) PM10max 22.4 27.55 40.89 22 29.73 43.5 0.4147 (>0.05) NOmax 11.07 35.66 82.02 11.1 28.43 77.36 0.9342 (>0.05) NO2max 40.21 54.79 67.28 40.12 53.04 68.63 0.8793 (>0.05) NOXmax 55.79 103.91 175.8 55.86 91.64 181.47 0.8761 (>0.05)
@@
Table 7 presented the quartile cutoff values of key clinical and environmental factors along with t-test p-values, significant differences between STEMI and NSTEMI cases emerge, highlighting distinct chronic and acute influences. Systolic BP is significantly higher in NSTEMI (75th percentile = 156) compared to STEMI (75th percentile = 146, p = 0.00187), suggesting that chronic hypertension is more prevalent in NSTEMI, whereas lower BP in STEMI reflects acute cardiovascular dysfunction. BMI follows a similar trend, with higher values in NSTEMI (75th percentile = 32.23 vs. 29.5 in STEMI, p = 0.00207), reinforcing the role of chronic obesity. Creatinine levels are also significantly elevated in STEMI (p = 0.00490), indicating potential acute kidney injury or pre-existing renal impairment as contributors. Notably, PM10 exposure is significantly higher in STEMI cases (p = 0.04867), suggesting that short-term environmental triggers may precipitate acute coronary events, while NO, NOX, and NO2 levels showed no significant difference (p $>$ 0.05), implying a lesser role in STEMI/NSTEMI differentiation. However, the lack of statistical significance in several AQIs may be due to the fact that STEMI and NSTEMI cases occurred on the same day, reducing the distinction between pollution-related effects. Age at admission trends slightly higher in NSTEMI (75th percentile = 81 vs. 80.25 in STEMI), but the difference is not statistically significant (p = 0.49108). These findings emphasize the chronic nature of NSTEMI, driven by hypertension and obesity, versus the acute nature of STEMI, potentially triggered by air pollution exposure, highlighting the need for targeted prevention strategies such as improved hypertension and obesity management and air quality interventions to mitigate ACS risk.
@@ 6 Discussion
Traditional cardiovascular risk scores, such as QRISK3, FRS, and GRACE, primarily rely on clinical data and do not include dynamic environmental influences. By incorporating air pollution data, the TabulaTime model addresses this gap, providing a more comprehensive risk assessment tool. This integration is particularly relevant given the growing body of evidence linking air pollution to adverse cardiovascular outcomes. For instance, pollutants such as PM10, NO2, NO, and NOx have been shown to exacerbate cardiovascular conditions through mechanisms like inflammation and oxidative stress [10, 63]. In this work, the presence of environmental factors like PM10 and NO2 among the more important features indicates that air quality and pollution may be significant predictors in the context of the model, possibly relating to respiratory or cardiovascular conditions. Our experimental results demonstrate that including air pollution data improves the accuracy of ACS risk prediction by 10.1%, a substantial gain over traditional clinical-only models. This finding aligns with prior epidemiological studies linking air pollutants (PM10, NO2, NOx) with increased cardiovascular stress, inflammation, and thrombotic events. By incorporating real-time exposure data, TabulaTime provides a more holistic assessment of ACS risk, highlighting the immediate effects of environmental exposure in triggering cardiovascular events.
One of the major innovations in TabulaTime is the PatchRWKV module, an efficient time-series feature extraction module that combines Recurrent Neural Networks (RNNs) and Transformer-based self-attention mechanisms. Our experiments across multiple classification and forecasting tasks confirmed that PatchRWKV surpassed state-of-the-art models such as PatchTST, Autoformer, and TimesNet, while maintaining linear computational complexity. The patching strategy, which isolates local temporal patterns (e.g., daily pollution cycles), allows the model to discern subtle exposure-response relationships that manual feature engineering might miss. This capability is critical for public health applications, where identifying high-risk periods (e.g., days with PM10 ¿40 µg/m³) can inform real-time interventions.

### v2

@@ 4.3.1 Classification Metrics
ROC Curve and AUC (Area Under the Curve): The ROC curve is a graphical representation of a classification model’s performance across all classification thresholds. It plots the True Positive Rate (TPR) against the False Positive Rate (FPR). The AUC provides a single scalar value to summarise the model’s performance; a higher AUC indicates better model performance.
@@ 4.3.2 Forecasting Metrics
4.4 Experimental configuration
4.4.1 Data Preprocessing
Data preprocessing is a critical step before model training. For the clinical tabular data from the MINAP dataset, several steps were taken. Categorical variables were converted into numerical format using One-Hot encoding. Numerical features were standardised to have a mean of 0 and a standard deviation of 1 to ensure consistency in data scale and improve model convergence.
Handling missing values is essential for robust model performance. In our dataset, missing values were relatively sparse, constituting approximately 2% of the clinical data. Our primary analysis employed K-Nearest Neighbours (KNN) imputation. This method estimates missing values by referencing the five most similar complete records (’neighbour’) based on other available features. KNN imputation is advantageous as it preserves local data structure and inter-feature relationships, making it particularly suitable for datasets with a low proportion of missing data like ours.
To justify the selection of KNN, we conducted a sensitivity analysis comparing three common imputation strategies:
  1. 1.
    Mean imputation: Replacing missing values with the mean of the respective feature.
  2. 2.
    K-Nearest Neighbours (KNN) imputation (k=5): The primary method used in this study.
  3. 3.
    Multivariate Imputation by Chained Equations (MICE): An iterative regression-based method that models each variable with missing values as a function of the other variables.
The comparative performance metrics (Accuracy and AUC) for the TabulaTime model using data imputed by these three methods are summarised in Table 2.
Imputation Method Accuracy AUC Mean Imputation 0.79 0.87 MICE 0.81 0.89 KNN Imputation (k=5) 0.82 0.91
Table 2: A sensitivity analysis comparing three primary imputation methods was conducted to support the choice of KNN.
These results indicate that KNN imputation achieved the best overall model performance in terms of both Accuracy and AUC, with MICE yielding comparable results. Both KNN and MICE significantly outperformed simple mean imputation. This sensitivity analysis confirms that the model’s performance is robust across different sophisticated imputation techniques and supports our decision to adopt KNN imputation as the primary method for handling missing data in this study.
4.4.2 Implementation Details
All experiments were implemented using Python 3.9. The deep learning models, including TabulaTime and its PatchRWKV module, were built using the PyTorch (version 2.5). Traditional machine learning models (Random Forest, LightGBM, CatBoost) were implemented using the Scikit-learn library. The source code for the TabulaTime framework and the experiments conducted in this study is publicly available at: https://gitlab.com/han-research/tabulatime.
Embedding Dimensions: The specific dimensions used for the input embeddings are as follows:
  • Tabular embedding ($X_{tab}$): The pre-processed clinical tabular data (including 21 clinical features and 10 air pollution statistical features) is embedded into a dense vector of dimension D = 31 using an MLP.
  • Patch embedding ($X_{ts}$): Hourly environmental data is segmented into non-overlapping daily patches (Patch size P = 24, Stride S = 24). Each patch is flattened and linearly projected into an embedding space with dimension C = 128. The resulting tensor has shape $X_{ts}\in\mathbb{R}^{N\times T\times C}$, where N is the number of environmental indicators (channels) and T is the number of patches (days, e.g., 10 for a 10-day sequence).
Training Configuration:
For model training and evaluation, the data was partitioned into training (80%) and testing (20%) subsets, ensuring temporal order was maintained where applicable. A validation subset (10% of the training data) was used for hyperparameter tuning and early stopping. The TabulaTime model, including the PatchRWKV module (configured with 4 layers and an embedding dimension C=128), was trained using the Adam optimiser with a learning rate of 0.001 and a batch size of 16. The training process was run for a maximum of 100 epochs, employing early stopping with a patience of 10 epochs based on the validation loss (Cross-Entropy Loss for classification tasks) to prevent overfitting. Baseline machine learning models (RF, LightGBM, CatBoost) were trained using their default hyperparameters as provided by Scikit-learn, representing standard configurations.
@@ 5.1 The Performance Evaluation of TabulaTime in ACS Prediction
The evaluation of the TabulaTime model was conducted through a series of experiments to assess its predictive performance for description of Acute Coronary Syndrome (ACS) risk. The experiments were designed to evaluate the influence of time-series air pollution data, compare TabulaTime with existing machine learning models, and analyse its generalisation capabilities under varying conditions.
@@ 5.1.1 Impact of Time-Series Length for ACS Prediction
The model was tested with input sequences corresponding to varying preceding durations (3, 7, 10, and 14 days). For the optimal 10-day window, this involved processing 10 consecutive 24-hour patches derived from the hourly air pollution data. As shown in Table 3, the 10-day time series achieved the best results, with an accuracy of 82.3% and an AUC of 0.91, outperforming other time lengths across all metrics. This suggests that a 10-day window strikes a balance between capturing essential trends and avoiding noise introduced by excessive data. Shorter windows (3 and 7 days) lacked critical information, while longer windows (14 days) diluted key patterns.
@@ 5.1.2 Comparison With Existing Machine Learning Models for ACS Risk Prediction
Table 4: Comparison of classification accuracy (%) for STEMI vs. NSTEMI prediction using traditional machine learning models (Random Forest, LightGBM, CatBoost) and the proposed TabulaTime method, under scenarios with and without air pollution data.
@@
Both models showed an improvement when air pollution is included as a factor. The performance of RF improves by 4.5% (from 0.6 to 0.627), LightGBM improved by 1.6% (from 0.645 to 0.655) and Catboost improves by 3.5% % (from 0.665 to 0.688). TabulaTime's performance improved by 10.1% (from 0.753 to 0.829). This indicated that incorporating air pollution data enhanced the predictive capabilities of both models, but the improvement was more pronounced for TabulaTime. Figure 5 reported the Area Under the Receiver Operating Characteristic Curve (AUC-ROC) for the best-performing machine learning model (Catboost) and the proposed TabulaTime. The results confirm that the model performance improved with the inclusion of air pollution indicators. The ROC of the Catboost increased from 0.67 to 0.72, while the proposed TabulaTime improved the prediction accuracy from 0.83 to 0.91.
Figure 5: ROC curve of Catboost(a) and TabulaTime(b) with and without air pollution.
@@ 5.2.1 Classification
Table 5: Classification accuracy (%) of various deep learning models on multivariate medical time series datasets including heartbeat monitoring, self-regulation of slow cortical potentials (SCP), and acute coronary syndrome (ACS) prediction.
@@ 5.2.2 Forecasting
Table 6: Performance comparison of various deep learning models on time series forecasting tasks for weather and air pollution using MSE and MAE metrics.
Figure 7: The forecasting result of PatchRWKV on Weather (a and b) and Salford Air Pollution dataset (c,d,e and f).
Table 6 presents the performance of various time series processing model on two forecasting tasks. The models are categorised into four types: MLP-based, CNN-based, Transformer-based, and RNN-based. Performance was evaluated using two metrics: Mean Squared Error (MSE) and Mean Absolute Error (MAE). The PatchRWKV approach demonstrates superior performance on both datasets, achieving the lowest MSE and highly competitive MAE values. Figure 7 showed the forecasting results of PatchRWKV on the two real datasets forecasting tasks. The model results were successful in predicting the trends of the pollution indicators. This indicates that PatchRWKV is highly effective in terms of both accuracy and robustness compared to other methods evaluated. Meanwhile, the RNN-based models generally outperformed other types in weather forecasting, with PatchRWKV and LSTM showing particularly strong results.
@@ 5.3 Feature Importance and Model Interpretability
Figure 8: Feature importance for ACS subtype prediction (STEMI vs. NSTEMI). Features include clinical, environmental, and historical variables.
In this work, we reported feature importance to explain how models make decisions to predict STEMI VS. NSTEMI.The analysis highlights the contributions of 31 risk factors (Figure 8), encompassing patient demographics, clinical history, vital signs, medications, diagnostic indicators, and environmental factors.
The analysis revealed that Systolic Blood Pressure (SBP) was the most influential predictor (importance score: 9.45), followed closely by Symptom-to-Admission Time (8.29) and Body Mass Index (BMI) (7.69), emphasising the critical roles of cardiovascular function, presentation timing, and metabolic factors. Environmental and physiological markers, such as average temperature (Tempavg) (6.13) and Creatinine levels (5.97), along with air quality metrics like maximum PM10 (PM10max) (5.37), were also significant contributors. Moderately important features included Age at Admission (5.26), average PM10 (PM10avg) (4.91), minimum temperature (Tempmin) (4.06), maximum NO (NOmax) (4.06), Smoking Status (3.92), and Previous Angina (3.85). Factors like Diabetes (2.98), Statin use (2.97), and Hypertension (1.55) showed relevance but had lower importance scores, suggesting a more indirect role in differentiating STEMI vs. NSTEMI in this model.
Figure 9: Feature Importance by Day for Air Pollution Exposure in ACS Prediction.
@@
Figure 9 showed the average importance of air quality each day for predicting ACS risk. The daily step importance analysis revealed that pollution levels on the second day prior to clinical presentation (Day 2) held the highest predictive significance (importance score: 0.0285), closely followed by the day immediately prior (Day 1, importance score: 0.0228) and the third day prior (Day 3, importance score: 0.0223). There was a general trend of declining relevance for days further preceding the event, with Day 9 having the lowest importance (0.0119). Notably, Day 10 exhibited slightly higher importance (0.0176) than Day 9. This finding highlights the acute impact of short-term pollution exposure on the presentation of ACS, while also suggesting potential influences from earlier exposures or possible modelling artefacts.
@@
STEMI NSTEMI T-test p-value 25% 50% 75% 25% 50% 75% Systolic BP 111.25 127.50 146.00 120.00 136.50 154.00 0.0019 (<0.05) SymptomToAdmissionTime 1.02 5.37 20.33 -213.04 6.03 33.74 0.0020 (<0.05) BMI 23.32 26.04 29.89 24.52 28.03 32.23 0.0021 (<0.05) Tempavg 7.60 11.26 15.17 7.12 10.61 15.03 0.0039 (<0.05) Creatinine 72.25 87.00 100.75 72.00 87.00 107.00 0.0049 (<0.05) PM10max 22.33 28.54 41.25 22.42 30.65 44.86 0.0487 (<0.05) Age At Admission 59.25 71.00 81.00 59.00 72.00 81.00 0.4911 (>0.05) PM10avg 10.61 12.94 17.04 10.62 13.68 19.24 0.0021 (<0.05) Tempmin 2.20 6.20 9.20 2.20 5.70 9.20 0.5147 (>0.05) NOavg 2.44 4.73 10.01 2.35 4.23 10.29 0.0487 (<0.05) NOmax 11.86 33.46 80.17 12.02 29.70 87.35 0.9342 (>0.05) NO2max 39.67 54.10 66.38 41.09 53.75 69.48 0.8793 (>0.05) NOXavg 19.31 28.95 43.00 19.67 27.09 43.95 0.628 (>0.05) NO2avg 15.47 20.80 28.03 15.63 20.54 28.49 0.4883 (>0.05) NOXmax 56.09 100.14 177.60 57.92 94.67 196.00 0.8761 (>0.05)
@@
Table 8 summarizes quartile values and t-test results comparing key predictors between STEMI and NSTEMI cases. Significant differences (p <0.05) were observed for several factors. NSTEMI patients showed significantly higher Systolic Blood Pressure, BMI, Creatinine levels, maximum PM10, average PM10, and average NO compared to STEMI patients. Symptom-to-Admission Time and average temperature also differed significantly between the groups.
Conversely, no significant differences (p >0.05) were found for Age at Admission, minimum temperature, or several nitrogen oxide metrics (NOmax, NO2max, NOXavg, NO2avg, NOXmax). These findings suggest NSTEMI presentation in this cohort is associated more strongly with markers of chronic conditions (hypertension, obesity), elevated creatinine, and higher exposure to PM10 and average NO.
@@ 6 Discussion
6.1 Model Performance, Innovations, and Clinical Insights
Traditional cardiovascular risk scores, such as QRISK3, FRS, and GRACE, primarily rely on clinical data and often ignore environmental influences like air pollution. TabulaTime addresses this gap by integrating time-series air pollution data with clinical factors, providing a more comprehensive risk assessment framework. This integration is crucial, given the substantial evidence linking pollutants such as PM10, NO2, NO, and NOx to adverse cardiovascular outcomes via mechanisms like inflammation and oxidative stress [10, 66]. Our feature importance analysis (Figure 8) supports this, identifying air quality metrics, specifically PM10 (maximum and average) and average NO, as significant contributors alongside key clinical factors in differentiating ACS presentations. Crucially, our experiments demonstrate that incorporating air pollution data significantly enhances predictive performance, improving the accuracy of TabulaTime by 10.1% compared to using clinical data alone. This finding aligns with epidemiological studies linking short-term air pollutant exposure (including PM10 and NOx) with increased cardiovascular events. By leveraging real-time environmental exposure data, TabulaTime offers a more holistic and timely assessment of ACS risk, capturing the immediate impact of environmental triggers.
One of the major innovations in TabulaTime is the PatchRWKV module, an efficient time-series feature extraction module that combines Recurrent Neural Networks (RNNs) and Transformer-based self-attention mechanisms. Our experiments across multiple classification and forecasting tasks confirmed that PatchRWKV surpassed state-of-the-art models such as PatchTST, Autoformer, and TimesNet, while maintaining linear computational complexity. The patching strategy, which isolates local temporal patterns (e.g., daily pollution cycles), allows the model to discern subtle exposure-response relationships that manual feature engineering might miss. This capability is critical for public health applications, where identifying high-risk periods (e.g., days with PM10 >40 µg/m³) can inform real-time interventions.